OpenStudy (anonymous):

Which operation is performed immediately before multiplication when simplifying the expression 18 divided by the quantity 2 to the second power plus 5 end quantity times 3 minus 2?

OpenStudy (danjs):

\[\frac{ 18 }{ 2^2 +5}*3-2\]

OpenStudy (anonymous):

that equation looks familiar

OpenStudy (anonymous):

Well that's the equation of the word problem u stated above

OpenStudy (anonymous):

no i asked this qwestion yesterday thats why but my computer crashed and i never got to use it

OpenStudy (anonymous):

ohhhh using BEDMAS B=Brackets E=Exponents M/D - multiplication and division A/S =addition or subtraction So which of the these operations are performed first

OpenStudy (danjs):

square the 2 first

OpenStudy (danjs):

\[\frac{ 18 }{ 2^2 +5}*3-2 = \frac{ 18 }{ 4+5 }*3 - 2 = \frac{ 18*3 }{ 9 } - 2 = \frac{ 54 }{ 9 }-2\]

OpenStudy (danjs):

turn the 2 into a fraction over 9 by multiplying 2 by 9/9 \[\frac{ 54 }{ 9 }-\frac{ 18 }{ 9 } = \frac{ 54-18 }{ 9 }\]